3. Let X be random variable with probability density function x(x)4 for 0 x 1, (Note: fx (x) = 0 outside this domain.) (a) Find E[X] and Var[X] (b) Let Y- X2 +5. Find E[Y] and Var[Y]. (c) Find PX 112 ).
Answer #1

3.

(a)

1 1 .1 0

3 5 1-5 4

Var[X] = E[X^2] - ( E[X])^2

1 .1 1

=4left [ rac{1^6}{6}-rac{0^6}{6} ight ] = rac{4}{6}

Var[X] = rac{4}{6}- left ( rac{4}{5} ight )^2 =var(X) ÷ (:). 00267 Var[X]

(b)

We apply transformation :

Y = X^2 + 5

=> X = sqrt{Y-5}

Jacobian = |J| =left | rac{mathrm{d} x }{mathrm{d} y} ight | = left | rac{mathrm{d} sqrt{y-5} }{mathrm{d} y} ight | = rac{1}{2}(y-5)^{1/2-1}

= rac{1}{2}(y-5)^{-1/2}

So, the density function of Y will be :

f(y) = 4x^3.|J| = 4(sqrt{y-5})^3. rac{1}{2}(y-5)^{-1/2} = 2(y-5)^{3/2-1/2} = 2(y-5)

Since,

0leq xleq 1

=>0leq sqrt{y-5}leq 1

=>5leq yleq 6

E[Y]=int_{5}^{6}yf(y)dy =int_{5}^{6}y (2(y-5))dy =int_{5}^{6}2(y^2-5y)dy =2left [ rac{y^3}{3}- 5rac{y^2}{2} ight ]_5^6

=2left [ rac{6^3}{3}- 5rac{6^2}{2}-left ( rac{5^3}{3}- 5rac{5^2}{2} ight ) ight ]

=5.667

Var[Y] = E[Y^2] - ( E[Y])^2

E[Y^2]=int_{5}^{6}y^2f(y)dy =int_{5}^{6}y^2 (2(y-5))dy =int_{5}^{6}2(y^3-5y^2)dy =2left [ rac{y^4}{4}- 5rac{y^3}{3} ight ]_5^6

=2left [ rac{6^4}{4}- 5rac{6^3}{3}-left ( rac{5^4}{4}- 5rac{5^3}{3} ight ) ight ]

=32.1667

Var[Y] = E[Y^2] - ( E[Y])^2 = 32.1667 - (5.667)^2 = 0.0518

(c)

P(X 2 1/3)

=rac{P(1/3leq Xleq 2/3)}{P(Xgeq 1/3)}

P(1/3leq Xleq 2/3) = int_{1/3}^{2/3}f(x)dx = int_{1/3}^{2/3}4x^3dx = 4left [ rac{x^4}{4} ight ]_{1/3}^{2/3}

= 4left [ rac{(2/3)^4}{4}-rac{(1/3)^4}{4} ight ]

0.185

P(Xgeq 1/3) =int_{1/3}^{1}f(x)dx = int_{1/3}^{1}4x^3dx = 4left [ rac{x^4}{4} ight ]_{1/3}^{1}

= 4left [ rac{(1)^4}{4}-rac{(1/3)^4}{4} ight ]

=0.988

P(X 〈 2/3.X-1/3) _ 0.185 0.1875 P(X < 2/3X 1/3)